OpenStudy (anonymous):

Let R be the region in the first quadrant enclosed by the grpah of y=√(6X-4), the line y=2X and the y-axis. a) Find the area of R. b) Set up, but do not integrate, an integral expression in terms of a single variable for the volume of the solid generated when R is revolved about the x-axis c) Set up, but do not integrate, an integral expression in terms of a single variable for the volume of the solid generated when R is revolved about the y-axis

sam (.sam.):

The line and the curve intersect at x=0 and x=2 Integrate sqrt(6x+4) - 2x between these two limits ∫ ((6x+4)^(1/2) - 2x) dx , from 0 to 2 let 6x+4 = u 6 dx = du dx =(1/6) du 6x=u-4 x=-(u-4)/6 The integral becomes (1/6) ∫ u^(1/2) du -(2/6)(1/6) ∫ (u-4) du (1/6)∫ u^(1/2) du -(1/18) ∫ (u-4) du ----(1) 6x+4=u when x=0, u=4 when x=2, u=16 Integrating (1) (1/6)(2/3) u^(3/2) from 4 to 16 -(1/18) u^2/2 from 4 to 16 + (4/18) u from 4 to 16 =6.2222-6.66667+2.66666 =2.2222